OpenStudy (anonymous):

find the derivative 6(square root)x

OpenStudy (inkyvoyd):

d/dx 6sqrtx

OpenStudy (inkyvoyd):

=d/dx 6*x^(1/2)

OpenStudy (inkyvoyd):

Remember, d/dx x^n = nx^(n-1), for n=/=1.

OpenStudy (inkyvoyd):

@bluemustangjosh , tell me what you got.

OpenStudy (anonymous):

i got the same thing thanks i just wanted to be sure

OpenStudy (inkyvoyd):

No, tell me what you got :D

OpenStudy (inkyvoyd):

*and sorry, n can be equal to 1

OpenStudy (anonymous):

well thats what it is isnt it you multiply the six by 1/2 and then 1/2 turns negative cause of the substraction so 3x^-1/2...

OpenStudy (anonymous):

am i right lol....

OpenStudy (inkyvoyd):

Yes. Make sure you put -1/2 in parenthesis if you type it.

OpenStudy (anonymous):

thanks a bunch

OpenStudy (inkyvoyd):

No problem :)

OpenStudy (anonymous):

wait i have another question sorry lol so if i were to change it back to a square root would it be 3(square root)x

OpenStudy (inkyvoyd):

That's the part that confuses a lot of people. x^(1/2) =sqrt x.

OpenStudy (inkyvoyd):

they are two different ways of writing the same thing.